OpenStudy (anonymous):

Find the exact value of cos^(-1)(cos(17pi/5))?

OpenStudy (chrisdbest):

Can you put that in equation form using the equation button located on the left of the "Post" button?

OpenStudy (anonymous):

\[\cos^{-1}(\cos( 17\pi/5))\]

OpenStudy (chrisdbest):

Aye!

OpenStudy (chrisdbest):

Now that's better!

OpenStudy (chrisdbest):

Based on my calculations, \[\frac{ 3\pi }{ 5 }\]

OpenStudy (jdoe0001):

@_alex_urena_ what's the \(cos(\pi)?\)

OpenStudy (anonymous):

what do you mean @jdoe0001 ?

OpenStudy (anonymous):

how did you get that @chrisdbest ?

OpenStudy (jdoe0001):

ohh, just that... if you were to get the \(cos(\pi)\) what would that give you?

OpenStudy (chrisdbest):

Sure, I'll tell you

OpenStudy (anonymous):

-1? @jdoe0001

OpenStudy (jdoe0001):

ok... so... what is now the \(cos^{-1}(-1)?\)

OpenStudy (anonymous):

pi

OpenStudy (jdoe0001):

yeap.... thus... one sec

OpenStudy (jdoe0001):

\(\bf cos(\pi )={\color{brown}{ -1}}\qquad cos^{-1}({\color{brown}{ -1}})=\pi \\ \quad \\ \textit{thus we could say that }cos^{-1}[cos(\pi )]=\pi \\ \quad \\ \textit{thus, we could also say that }cos^{-1}\left[ cos\left( \cfrac{17\pi }{5} \right) \right]\implies \cfrac{17\pi }{5}\)

OpenStudy (chrisdbest):

What he's trying to say is that when you have an inverse cosine and a cosine, they cancel out

OpenStudy (jdoe0001):

so in short, \(\bf cos^{-1}[cos(whatever)]=whatever\qquad \\ \quad \\sin^{-1}[sin(whatever)]=whatever \\ \quad \\ tan^{-1}[tan(whatever)]=whatever\)

OpenStudy (chrisdbest):

Lols, that's basically what I said

OpenStudy (anonymous):

so the answer is 17pi/5 ?

OpenStudy (chrisdbest):

It might, I thought it was 3pi/5

OpenStudy (jdoe0001):

yeap, unless, the inverse function apply, which so far I don't see they do

OpenStudy (jdoe0001):

inverse functions restrictions I meant, doesn't seem like in this context they do

OpenStudy (chrisdbest):

Because the way you wrote it in equation form, I plugged that into my calculator the same way, and I got 3pi/5

OpenStudy (anonymous):

hmmm well how did you get 3pi/5 @chrisdbest

OpenStudy (chrisdbest):

Because the way you wrote it in equation form, I plugged that into my calculator the same way, and I got 3pi/5

OpenStudy (anonymous):

ok but how would you do without the calculator? or do you have to use one to solve this problem?

OpenStudy (chrisdbest):

I use one to solve the problem.

OpenStudy (anonymous):

hmmm ok thnx